# Moduli spaces of untwisted wild Riemann surfaces

### v1

@@ 4.1.2. Complex manifolds as sheaves
4.1.3. A moduli problem which is not a sheaf
[…]
4.1.4. Stacks
The problem in the above example comes from the following phenomenon: over each $B_{i}$ we have an isomorphism $f_{i}\colon\Sigma\rightarrow\Sigma_{0}$, but over $B_{1}\cap B_{2}$ the isomorphisms $f_{1},f_{2}$ differ by a non-trivial automorphism of $\Sigma_{0}$ (sending $y$ to $-y$). Hence, we see that the existence of non-trivial automorphisms of elliptic curves is intimately related to the fact that $M^{/iso}_{1,1}\colon\mathrm{Man}_{\mathbf{C}}^{op}\rightarrow\mathrm{Sets}$ is not a sheaf. Grothendieck’s idea to handle this issue is to make automorphisms part of the picture: rather than looking at functors $M\colon\mathrm{Man}_{\mathbf{C}}^{op}\rightarrow\mathrm{Sets}$, one considers (2-)functors
@@ 4.1.4. Stacks
We say that $\mathsf{M}\colon\mathrm{Man}_{\mathbf{C}}^{op}\rightarrow\mathrm{Groupoids}$ is a stack if it satisfies the following two conditions (cf. [Hei05, Definition 1.1, Remark 1.2]); below the pullback of an object $s\in\mathsf{M}(B)$ to an open $B^{\prime}\subset B$ is denoted by $s_{|B^{\prime}}$.
  1. (1)
    Objects glue: take an open cover $(B_{i})_{i\in I}$ of a manifold $B$, objects $s_{i}\in\mathsf{M}(B_{i})$ and isomorphims $\varphi_{ij}\colon s_{i|B_{i}\cap B_{j}}\rightarrow s_{j|B_{i}\cap B_{j}}$ satisfying the cocycle condition $\varphi_{jk}\circ\varphi_{ij}=\varphi_{ik}$ on triple intersections $B_{i}\cap B_{j}\cap B_{k}$. There exists $s\in\mathsf{M}(B)$ together with isomorphisms $\varphi_{i}\colon s_{|B_{i}}\rightarrow s_{i}$ such that $\varphi_{ij}=\varphi_{j}\circ\varphi_{i}^{-1}$.
  2. (2)
    Isomorphisms glue: take an open cover $(B_{i})_{i\in I}$ of a manifold $B$, and two objects $s,s^{\prime}\in\mathsf{M}(B)$. Given isomorphisms $\varphi_{i}\colon s_{|B_{i}}\rightarrow s^{\prime}_{|B_{i}}$ which agree on the intersections $B_{i}\cap B_{j}$, there is a unique $\varphi\colon s\rightarrow s^{\prime}$ restricting to $\varphi_{i}$ on each $B_{i}$.
@@
4.1.5 Remark.
There is a notion of morphism between two morphisms of stacks, which we will call transformation of morphisms. Namely, let $F_{1},F_{2}\colon\mathsf{M}\rightarrow\mathsf{N}$ be morphisms of stacks. By definition, for each $B$ we have functors $F_{1,B},F_{2,B}\colon\mathsf{M}(B)\rightarrow\mathsf{N}(B)$. A transformation from $F_{1}$ to $F_{2}$ is a collection of natural transformations from $F_{1,B}$ to $F_{2,B}$ for every complex manifold $B$, compatible with pullbacks (cf. [Hei05, Remark 3.1.4]; formally, stacks form a 2-category with invertible 2-morphisms).
4.1.6. Atlases and analytic stacks
@@
4.1.7 Remark.
@@
Let us point out that the above notion of representability is very strong: if $\mathsf{M},\mathsf{N}$ are (stacks attached to) manifolds then we are asking in particular that all the fibres of $F$ are manifolds, which is far from true for a general holomorphic map. A much better approach to the general theory consists in replacing $\mathrm{Man}_{\mathbf{C}}$ by the category of complex analytic spaces, which admits fibre products (cf. also [BN06, Remark 3.4]). However, the morphisms of stacks of interest in this document will satisfy the above condition. Hence, to make the text more accessible, we have chosen to work with the above definition.
4.1.8 Definition.
@@
A stack $\mathsf{M}$ is called analytic stack if there is a manifold $X$ and a morphism of stacks $\underline{X}\rightarrow\mathsf{M}$ which is a surjective submersion, i.e. a representable map such that for every manifold $B$ and every morphism $\underline{B}\rightarrow\mathsf{M}$ the projection $\underline{B}\times_{\mathsf{M}}\underline{X}\rightarrow\underline{B}$ is (induced by) a surjective holomorphic submersion of complex manifolds. Any manifold $\underline{X}\rightarrow\mathsf{M}$ with the previous property is called an atlas of the stack $\mathsf{M}$.
4.1.9. Moduli stacks of curves with marked points
@@
Let $g,m\geq 0$ be two nonnegative integers. The (2)-functor $\mathsf{M}_{g,m}$ sending a complex manifold $B$ to the groupoid of families of compact Riemann surfaces of genus $g$ over $B$ with $m$ (ordered) disjoint families of marked points is an analytic stack. More precisely, it is a quotient stack $[T_{g,m}/\Gamma_{g,m}]$, where $T_{g.m}$ is a Teichmüller space. This follows from the description of the functor of points of $T_{g,m}$ as a moduli space of Riemann surfaces with marked points and “Teichmüller structure” (cf. [Gro62, Théorème 3.1] for the case of $m=0$).
4.2. The fundamental group of an analytic stack
@@
In this section we give the definition of the (topological) fundamental group of an analytic stack, borrowed from [Noo05]. This is done in two steps: firstly, one defines the topological stack “underlying” an analytic stack; secondly, one extends the usual definition of fundamental group of a topological space to topological stacks. The first step rests on the dictionary between analytic stacks and (complex analytic) groupoids, which gives a concrete way to represent stacks in terms of “manifolds generators and relations”.
4.2.1. Presentation of analytic stacks via groupoids
@@
Let $\mathsf{M}$ be an analytic stack and $\underline{X}\rightarrow\mathsf{M}$ an atlas. The fibre product $\underline{X}\times_{\mathsf{M}}\underline{X}$ is of the form $\underline{R}$ for a manifold $R$ endowed with two projections to $X$, which define the source and target morphism of a groupoid $X_{\bullet}$ (in $\mathrm{Man}_{\mathbf{C}}$). Conversely, each groupoid $R\rightrightarrows X$ gives rise to a quotient stack $[X/R]$ [Noo05, §3.2], [Hei05, §3]. Heuristically, the groupoid $X_{\bullet}$ can be thought of as a “presentation” of the stack $\mathsf{M}$: the two projection maps describe which points of $X$ should be identified to obtain $\mathsf{M}$. However, it can happen that a point $x\in X$ gets identified with itself via more than one $r\in R$; this information is remembered by the stack $\mathsf{M}$ and is related precisely to non-trivial automorphisms of the objects parametrised by $\mathsf{M}$ but it would be lost looking at the quotient topological space $X/R$.
4.2.2. Topological stack associated with an analytic stack
@@
Let $\mathsf{M}$ be an analytic stack; choose an atlas $\underline{X}\rightarrow\mathsf{M}$ and let $X_{\bullet}=R\rightrightarrows X$ be the resulting groupoid. Consider the topological spaces $X^{top}$ and $R^{top}$ underlying $X$ and $R$, and the groupoid in topological spaces $X_{\bullet}^{top}=R^{top}\rightrightarrows X^{top}$; the associated stack on the category of topological spaces is denoted by $\mathsf{M}^{top}$, and is called the topological stack attached to $\mathsf{M}$ (note that $\mathsf{M}^{top}$ is a topological stack in the sense of [Noo05, Definition 13.8], using as local fibrations those coming from [Noo05, Example 13.1.3]). The fact that $\mathsf{M}^{top}$ is well defined, i.e. it does not depend on the choice of the atlas, follows from the next lemma.
4.2.3 Lemma.
@@
Let $\mathsf{M}$ be an analytic stack and let $\underline{X}_{1}\rightarrow\mathsf{M},\underline{X}_{2}\rightarrow\mathsf{M}$ be two atlases. The quotient stacks $[X^{top}_{1}/R^{top}_{1}]$ and $[X^{top}_{2}/R^{top}_{2}]$ are equivalent.
@@
The lemma can be proved looking at an atlas refining $X_{1},X_{2}$, as sketched in [Noo05, p. 79]. For completeness, let us give the details. The fibre product $\underline{X}_{1}\times_{\mathsf{M}}\underline{X}_{2}$ is of the form $\underline{X}_{3}$ for a manifold $X_{3}$. The projection $X_{3}\rightarrow X_{1}$ (resp. $X_{3}\rightarrow X_{2}$) is a (locally trivial) $X_{2,\bullet}$-bundle (resp. $X_{1,\bullet}$-bundle) in the sense of [Hei05, §3], and the two actions defining the bundle structures commute. Hence, the same properties are true for the maps $X_{3}^{top}\rightarrow X_{i}^{top}$ for $i=1,2$, and the lemma follows from [Hei05, Lemma 3.2]. ∎
4.2.4. The fundamental group
A pointed analytic stack is a couple $(\mathsf{M},x)$ consisting of an analytic stack $\mathsf{M}$ and a map ¯$\underline{*}\rightarrow\mathsf{M}$, corresponding to an object $x\in\mathsf{M}(*)$ (the image of the identity morphism of *). We denote by $(\mathsf{M}^{top},x)$ the associated pointed topological stack. Any pointed topological space $(T,x)$ gives rise to a pointed topological stack $(\underline{T},x)$. The following definition (coming from [Noo05, Definition 17.5]) is analogous to the usual definition of the fundamental group of a topological space; the subtlety and interest of the definition stem from the fact that the expressions “homotopy” and “map of pairs” have a more refined meaning in the world of stacks than in the context of topological spaces. This will be recalled below.
4.2.5 Definition.
The fundamental group of a pointed analytic stack $(\mathsf{M},x)$, denoted by $\pi_{1}(\mathsf{M},x)$, is the set of homotopy classes of maps of pairs $(\underline{S}^{1},1)\rightarrow(\mathsf{M}^{top},x)$, endowed with the group structure defined in [Noo05, pp. 59, 60].
4.2.6 Remark.
Let us give some comments on the previous definition. The main point is that the notions of map of pairs and homotopy between continuous maps of topological spaces can be expressed via the commutativity of certain diagrams. However, as explained in Remark 4.1.5, in the world of stacks we have a notion of transformation between morphisms; to extend the classical definitions to the context of stacks, one includes a transformation from a map $F_{1}$ to a map $F_{2}$ whenever usually one would have asked the two maps to be equal.

### v2

@@ A.1.2. Complex manifolds as sheaves
A.1.3. A moduli problem which is not a sheaf
Many moduli problems of interest turn out not to be representable, often because they are not sheaves on $\mathrm{Man}_{\mathbf{C}}$.
For example, consider the functor $M^{/iso}_{1,1}$ sending $B$ to the set of isomorphism classes of elliptic curves over $B$. These are couples ($\pi\colon\Sigma\rightarrow B,\sigma$) consisting of a proper submersion $\pi$ of complex manifolds whose fibres are one dimensional complex tori, and a holomorphic section $\sigma\colon B\rightarrow\Sigma$ of $\pi$; isomorphisms are biholomorphisms commuting with the structure map to $B$ and with the sections.
We claim that $M^{/iso}_{1,1}$ is not a sheaf on $\mathrm{Man}_{\mathbf{C}}$. Indeed, take $B=\mathbf{C}\smallsetminus\{0\}$ and consider the product $B\times\mathbf{P}^{2}(\mathbf{C})$, whose elements we denote by $(b,[x:y:z])$. Let $\Sigma\subset B\times\mathbf{P}^{2}(\mathbf{C})$ be the submanifold cut out by the equation $by^{2}z=x^{3}-z^{2}x$. Let $\pi\colon\Sigma\rightarrow B$ be the map induced by projection of $B\times\mathbf{P}^{2}(\mathbf{C})$ on the first component, and $\sigma$ be the map sending $b$ to $(b,[0:1:0])$. We can write $B=B_{1}\cup B_{2}$, where $B_{1}$ (resp. $B_{2}$) is the complement in $\mathbf{C}$ of the half-line $\mathbf{R}_{\geq 0}$ (resp. $\mathbf{R}_{\leq 0}$). Over each $B_{i}$ we can choose a holomorphic square root function $r_{i}\colon B_{i}\rightarrow\mathbf{C}$. Then, the map sending $(b,[x:y:z])$ to $(b,[x:r_{i}(b)y:z])$ induces an isomorphism between $\Sigma$ and the elliptic curve $\Sigma_{0}$ with equation $y^{2}z=x^{3}-z^{2}x$ over $B_{i}$. However, there is no isomorphism between $\Sigma$ and $\Sigma_{0}$ over the whole BB—the monodromy representation attached to $\Sigma$ is non-trivial, cf. [Lvo19, Proposition 3.3]—hence $M^{/iso}_{1,1}$ is not a sheaf.
A.1.4. Stacks
The problem in the above example comes from the following phenomenon: over each $B_{i}$ we have an isomorphism $f_{i}\colon\Sigma\rightarrow\Sigma_{0}$, but over $B_{1}\cap B_{2}$ the isomorphisms $f_{1},f_{2}$ differ by a non-trivial automorphism of $\Sigma_{0}$ (sending $y$ to y-y). Hence, we see that the existence of non-trivial automorphisms of elliptic curves is intimately related to the fact that $M^{/iso}_{1,1}\colon\mathrm{Man}_{\mathbf{C}}^{op}\rightarrow\mathrm{Sets}$ is not a sheaf.
Grothendieck’s idea to handle this issue is to make automorphisms part of the picture: rather than looking at functors $M\colon\mathrm{Man}_{\mathbf{C}}^{op}\rightarrow\mathrm{Sets}$, one considers (2-)functors
@@ A.1.4. Stacks
We say that $\mathsf{M}\colon\mathrm{Man}_{\mathbf{C}}^{op}\rightarrow\mathrm{Groupoids}$ is a stack if it satisfies the following two conditions (cf. [Hei05, Definition 1.1, Remark 1.2]). [^1] Below the pullback of an object $s\in\mathsf{M}(B)$ to an open $B^{\prime}\subset B$ is denoted by $s_{|B^{\prime}}$.
Objects glue:
take an open cover $(B_{i})_{i\in I}$ of a manifold $B$, objects $s_{i}\in\mathsf{M}(B_{i})$ and isomorphisms $\varphi_{ij}\colon s_{i|B_{i}\cap B_{j}}\rightarrow s_{j|B_{i}\cap B_{j}}$ satisfying the cocycle condition $\varphi_{jk}\circ\varphi_{ij}=\varphi_{ik}$ on triple intersections $B_{i}\cap B_{j}\cap B_{k}$. Then there exists $s\in\mathsf{M}(B)$ together with isomorphisms $\varphi_{i}\colon s_{|B_{i}}\rightarrow s_{i}$ such that $\varphi_{ij}=\varphi_{j}\circ\varphi_{i}^{-1}$.
Isomorphisms glue:
take an open cover $(B_{i})_{i\in I}$ of a manifold $B$, and two objects $s,s^{\prime}\in\mathsf{M}(B)$. Then, given isomorphisms $\varphi_{i}\colon s_{|B_{i}}\rightarrow s^{\prime}_{|B_{i}}$ which agree on double intersections $B_{i}\cap B_{j}$, there exists a unique $\varphi\colon s\rightarrow s^{\prime}$ restricting to $\varphi_{i}$ on each $B_{i}$.
@@
A.1.5 Remark.
There is a notion of morphism between two morphisms of stacks, which we will call transformation of morphisms.
Namely, let $F_{1},F_{2}\colon\mathsf{M}\rightarrow\mathsf{N}$ be morphisms of stacks. By definition, for each $B$ we have functors $F_{1,B},F_{2,B}\colon\mathsf{M}(B)\rightarrow\mathsf{N}(B)$. A transformation from $F_{1}$ to $F_{2}$ is a collection of natural transformations from $F_{1,B}$ to $F_{2,B}$ for every complex manifold $B$, compatible with pullbacks (cf. [Hei05, Remark 3.1.4]; formally, stacks form a 2-category with invertible 2-morphisms).
A.1.6. Atlases and analytic stacks
@@
A.1.7 Remark.
Let us point out that the above notion of representability is very strong: if $\mathsf{M},\mathsf{N}$ are (stacks attached to) manifolds then we are asking in particular that all the fibres of $F$ are manifolds, which is far from true for a general holomorphic map.
@@
A better approach to the general theory consists in replacing $\mathrm{Man}_{\mathbf{C}}$ by the category of complex analytic spaces, which admits fibre products (cf. also [BN06, Remark 3.4]). However, the morphisms of stacks of interest in this document will satisfy the above condition. Hence, to make the text more accessible, we have chosen to work with the above definition.
A.1.8 Definition.
@@
A stack $\mathsf{M}$ is called analytic stack if there is a manifold $X$ and a morphism of stacks $\underline{X}\rightarrow\mathsf{M}$ which is a surjective submersion, i.e. a representable map such that for every manifold $B$ and every morphism $\underline{B}\rightarrow\mathsf{M}$ the projection $\underline{B}\times_{\mathsf{M}}\underline{X}\rightarrow\underline{B}$ is (induced by) a surjective holomorphic submersion of complex manifolds.
Any manifold $\underline{X}\rightarrow\mathsf{M}$ with the previous property is called an atlas of the stack $\mathsf{M}$.
A.1.9. Moduli stacks of curves with marked points
@@
Let $g,m$ be two nonnegative integers. The (2-)functor $\mathsf{M}_{g,m}$ sending a complex manifold $B$ to the groupoid of families of compact Riemann surfaces of genus $g$ over $B$ with $m$ (ordered) disjoint families of marked points is an analytic stack. If $g\geq 1,m\geq 1$ or $g=0$ and $m\geq 3$, it is a quotient stack $T_{g,m}/\Gamma_{g,m}]$, where $T_{g.m}$ is a Teichmüller space and $\Gamma_{g,m}$ is the (pure) mapping class group. This follows from the description of the functor of points of $T_{g,m}$ as a moduli space of Riemann surfaces with marked points and “Teichmüller structure”; cf. [Gro62, Théorème 3.1], surveyed in [AJP16], for the case $m=0$, and [Eng75, Theorem 3.6, Proposition 3.9] or [Ear73] for the general case.
A.2. The fundamental group of an analytic stack
@@
In this section we give the definition of the (topological) fundamental group of an analytic stack, borrowed from [Noo]. This is done in two steps: firstly, one defines the topological stack “underlying” an analytic stack; secondly, one extends the usual definition of fundamental group of a topological space to topological stacks. The first step rests on the dictionary between analytic stacks and (complex analytic) groupoids, which gives a concrete way to represent stacks in terms of “manifolds generators and relations”.
A.2.1. Presentation of analytic stacks via groupoids
@@
Let $\mathsf{M}$ be an analytic stack and $\underline{X}\rightarrow\mathsf{M}$ an atlas. The fibre product $\underline{X}\times_{\mathsf{M}}\underline{X}$ is of the form $\underline{R}$ for a manifold $R$ endowed with two projections to $X$, which define the source and target morphism of a groupoid $X_{\bullet}$ (in $\mathrm{Man}_{\mathbf{C}}$). Conversely, each groupoid $R\rightrightarrows X$ gives rise to a quotient stack $[X/R]$ [Noo, § 3.2], [Hei05, § 3]. Heuristically, the groupoid $X_{\bullet}$ can be thought of as a “presentation” of the stack $\mathsf{M}$: the two projection maps describe which points of $X$ should be identified to obtain $\mathsf{M}$. However, it can happen that a point $x\in X$ gets identified with itself via more than one $r\in R$; this information is remembered by the stack $\mathsf{M}$—and is related precisely to non-trivial automorphisms of the objects parametrised by $\mathsf{M}$—but it would be lost looking at the quotient topological space $X/R$.
A.2.2. Topological stack associated with an analytic stack
@@
Let $\mathsf{M}$ be an analytic stack; choose an atlas $\underline{X}\rightarrow\mathsf{M}$ and let $X_{\bullet}=R\rightrightarrows X$ be the resulting groupoid. Consider the topological spaces $X^{top}$ and $R^{top}$ underlying $X$ and $R$, and the groupoid in topological spaces $X_{\bullet}^{top}=R^{top}\rightrightarrows X^{top}$; the associated stack on the category of topological spaces is denoted by $\mathsf{M}^{top}$, and is called the topological stack attached to $\mathsf{M}$ (note that $\mathsf{M}^{top}$ is a topological stack in the sense of [Noo, Definition 13.8], using as local fibrations those coming from [Noo, Example 13.1.3]). The fact that $\mathsf{M}^{top}$ is well defined, i.e. it does not depend on the choice of the atlas, follows from the next lemma.
A.2.3 Lemma.
@@
Let $\mathsf{M}$ be an analytic stack and let $\underline{X}_{1}\rightarrow\mathsf{M}$, $\underline{X}_{2}\rightarrow\mathsf{M}$ be two atlases. The quotient stacks $X^{top}_{1}/R^{top}_{1}]$ and $X^{top}_{2}/R^{top}_{2}]$ are equivalent.
@@
The lemma can be proved looking at an atlas refining $\underline{X}_{1}$, $\underline{X}_{2}$, as sketched in [Noo, p. 79]. For completeness, let us give the details. The fibre product $\underline{X}_{1}\times_{\mathsf{M}}\underline{X}_{2}$ is of the form $\underline{X}_{3}$ for a manifold $X_{3}$. The projection $X_{3}\rightarrow X_{1}$ (resp. $X_{3}\rightarrow X_{2}$) is a (locally trivial) $X_{2,\bullet}$-bundle (resp. $X_{1,\bullet}$-bundle) in the sense of [Hei05, § 3], and the two actions defining the bundle structures commute. Hence, the same properties are true for the maps $X_{3}^{top}\rightarrow X_{i}^{top}$ for $i=1,2$, and the lemma follows from [Hei05, Lemma 3.2]. ∎
A.2.4. The fundamental group
A pointed analytic stack is a couple $(\mathsf{M},x)$ consisting of an analytic stack $\mathsf{M}$ and a map $\underline{*}\rightarrow\mathsf{M}$, corresponding to an object $x\in\mathsf{M}(*)$ (the image of the identity morphism of *). We denote by $(\mathsf{M}^{top},x)$ the associated pointed topological stack. Any pointed topological space $(T,x)$ gives rise to a pointed topological stack $(\underline{T},x)$. The following definition (coming from [Noo, Definition 17.5]) is analogous to the usual definition of the fundamental group of a topological space; the subtlety and interest of the definition stem from the fact that the expressions “homotopy” and “map of pairs” have a more refined meaning in the world of stacks than in the context of topological spaces. This will be recalled below.
A.2.5 Definition.
The fundamental group of a pointed analytic stack $(\mathsf{M},x)$, denoted by $\pi_{1}(\mathsf{M},x)$, is the set of homotopy classes of maps of pairs $(\underline{S}^{1},1)\rightarrow(\mathsf{M}^{top},x)$, endowed with the group structure defined in [Noo, pp. 59, 60].
A.2.6 Remark.
Let us give some comments on the previous definition. The main point is that the notions of map of pairs and homotopy between continuous maps of topological spaces can be expressed via the commutativity of certain diagrams. However, as explained in Remark A.1.5, in the world of stacks we have a notion of transformation between morphisms; to extend the classical definitions to the context of stacks, one includes a transformation from a map $F_{1}$ to a map $F_{2}$ whenever usually one would have asked the two maps to be equal.
For instance, a map $(\underline{S}^{1},1)\rightarrow(\mathsf{M}^{top},x)$ consists by definition [Noo, Definition 17.1] of the datum of a morphism $F\colon\underline{S}^{1}\rightarrow\mathsf{M}^{top}$ together with a transformation between the composition $\underline{*}\xrightarrow{1}\underline{S}^{1}\rightarrow\mathsf{M}^{top}$ and the map $\underline{*}\rightarrow\mathsf{M}^{top}$ corresponding to $x$. Such a transformation is induced by an isomorphism between $F(1)$ and $x$. In other words, we do not ask $F$ to map $1$ to $x$, but only to an object isomorphic to $x$, and we incorporate isomorphisms $F(1)\simeq x$ into our definition of “loop”. In particular, we see that automorphisms of objects of $\mathsf{M}(*)$ contribute to the fundamental group of the stack, in a way explained more precisely in [Noo, § 17]. Finally, the notion of homotopy between maps of pairs is defined in [Noo, Definition 17.2].
A.2.7 Remark.
  1. (1)
    If $\mathsf{M}=\underline{X}$ for a manifold $X$, then the transformations mentioned in the previous remark are always trivial, hence Definition A.2.5 recovers the usual fundamental group of $(X,x)$.
  2. (2)